OpenStudy (idealist10):

Find the curvature of r(t)=ti+(1/2)t^2*j+t^2*k.

OpenStudy (idealist10):

@hartnn @myininaya @thomaster @iambatman @Compassionate

OpenStudy (idealist10):

@Hero @Luigi0210

Parth (parthkohli):

\[k = \dfrac{|\vec r'(t) \times \vec r ''(t)|}{|r'(t)|^3}\]In your case, \(\vec r' (t) = \hat i + t\hat j + 2t \hat k\) and \(\vec r''(t) = \hat j + 2\hat k\).

OpenStudy (idealist10):

So I got r'(t)=<1, t, 2t> and r"(t)=<0, 1, 2>, now what do I do?

Parth (parthkohli):

Plug that into the formula.

OpenStudy (idealist10):

So r'(t)xr"(t)=<0, t, 4t>?

Parth (parthkohli):

Well, I haven't tried it yet, but if you say so.

OpenStudy (idealist10):

So what's \[\left| r'(t)*r"(t) \right|\]

OpenStudy (anonymous):

Just follow what the formula tells you, that's really what's to it

OpenStudy (idealist10):

That's what I did and got r'(t)xr"(t)=<0, t, 4t>.

Parth (parthkohli):

It's the magnitude of the vector.

OpenStudy (idealist10):

So how do I find the magnitude of the vector?

Parth (parthkohli):

How do you find the magnitude of vectors?

Parth (parthkohli):

Determinant, lol.

Parth (parthkohli):

Holy hell, I just noticed that Idealist didn't calculate the cross product either.

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Do you know how to do the cross product?

OpenStudy (idealist10):

No.

OpenStudy (anonymous):

You will have a 3 by 3 determinant

Parth (parthkohli):